# Bi-Level Online Provisioning and Scheduling with Switching Costs and Cross-Level Constraints

### v1

@@ Abstract
We study a bi-level online provisioning and scheduling problem motivated by network resource allocation, where provisioning decisions are made at a slow time scale while queue-/state-dependent scheduling is performed at a fast time scale. We model this two-time-scale interaction using an upper-level online convex optimization (OCO) problem and a lower-level constrained Markov decision process (CMDP). Existing OCO typically assumes stateless decisions and thus cannot capture MDP network dynamics such as queue evolution. Meanwhile, CMDP algorithms typically assume a fixed constraint threshold, whereas in provisioning-and-scheduling systems, the threshold varies with online budget decisions. To address these gaps, we study bi-level OCO-CMDP learning under switching costs (budget reprovisioning/system reconfiguration) and cross-level constraints that couple budgets to scheduling decisions. Our new algorithm solves this learning problem via several non-trivial developments, including a carefully designed dual feedback that returns the budget multiplier as sensitivity information for the upper-level update and a lower level that solves a budget-adaptive safe exploration problem via an extended occupancy-measure linear program. We establish near-optimal regret and high-probability satisfaction of the cross-level constraints.
@@ I Introduction
Modern networking systems, such as 5G/6G network and edge computing, must support heterogeneous services (e.g., video streaming, augmented reality (AR), virtual reality (VR), and autonomous driving) over shared wireless infrastructure [8]. A central operational challenge in these systems is online resource provisioning under uncertainty. In particular, a network operator must decide how much bandwidth/compute/energy to allocate to a slice over time, balancing quality of service (QoS) against operational expenditure (OpEx). Practical systems exhibit multiple coupled time scales. As illustrated in our motivating example online bandwidth provisioning with queue-aware scheduling (see Fig. 1 and Sec. III for details), the operator provisions an episode-level resource budget at a slow time scale (e.g., bandwidth reserved for a slice over seconds/minutes). This budget directly incurs OpEx and influences the QoS that can be achieved through scheduling. Within each episode, the scheduler performs state-dependent, packet-level actions at a fast time scale (e.g., allocating resource blocks every millisecond). These actions are constrained by the provisioned resource budget and aim to reduce latency and congestion, improving QoS [9, 6]. Moreover, changing the provisioning level across episodes is not free. It induces reconfiguration overheads (e.g., hardware retuning, virtualization/migration cost, and instability), which are naturally captured by switching costs [16, 7].
Online convex optimization (OCO) is a classical abstraction for sequential resource allocation and has been extensively studied under adversarial and stochastic settings [18, 16, 7]. Recent work incorporates switching costs and related operational constraints to penalize aggressive changes in decisions, which is attractive for network slicing. However, OCO with switching costs typically does not model endogenous Markovian state evolution nor provide Markovian-style safety under unknown dynamics. For example, the per-round loss depends on the current decision and does not capture the evolution of system states, while in networked systems, state variables such as queue lengths evolve endogenously. Consequently, provisioning decisions that appear cost-effective in the short term may cause persistent congestion and latency violations.
@@
Motivated by the above gaps, we propose a bi-level online optimization and learning framework that integrates online budget provisioning with state-aware safe scheduling. At the upper level, an OCO operator optimizes and provisions budgets $\{b_{k}\}_{k=1}^{K}$ across episodes, incurring a convex service cost and a quadratic switching cost. At the lower level, the system evolves as an episodic CMDP under varying thresholds. Given the optimized provisioning budget $b_{k}$, the agent must learn a safe policy via RL under unknown dynamics. The two levels are coupled through an episode-wise budget constraint and an episode-wise objective that aggregates cross-level costs.
@@ II-A Upper-Level Budget Provisioning
At the slow time scale, at the beginning of each episode $k\in[K]$, the agent selects an upper-level budget $b_{k}\in\mathcal{B}\triangleq[B_{0},T]$, which can represent the episode-level resource budget provisioned for a network slice (bandwidth, compute cycles, or energy) over seconds/minutes. Choosing $b_{k}$ incurs an episode-dependent convex service cost $f_{k}(b_{k})$, which is not revealed before the decision is made and is revealed at the end of episode $k$. As is typical in online convex optimization with switching costs [7, 4, 16], the service cost $f_{k}(\cdot)$ could change arbitrarily, and it is $\theta_{f}$-strongly convex and differentiable on $\mathcal{B}$ with bounded gradients $\|\nabla f_{k}(\cdot)\|\leq F$.
@@ II-B Lower-Level Stateful Scheduling
At each time $t$ of episode $k$, the agent observes the current state $s_{k,t}$ (e.g., queue backlogs and channel conditions), and then selects a (possibly randomized) scheduling/allocation action $a_{k,t}\sim\pi_{k,t}(\cdot\mid s_{k,t})$, where $\pi_{k}^{\text{A}}=(\pi_{k,t})_{t=1}^{T}$ is a policy with $\pi_{k,t}:\mathcal{S}\to\Delta(\mathcal{A})$. The agent then incurs a stage loss $l_{t}(s_{k,t},a_{k,t})$ (e.g., delay/backlog penalty), consumes resources $d_{t}(s_{k,t},a_{k,t})$ (e.g., used resource blocks/power), and the system evolves to next state $s_{k,t+1}\sim P_{t}(\cdot\mid s_{k,t},a_{k,t})$ due to unknown arrivals and service dynamics.
@@
$V_{r}(\pi;P)$ $\triangleq\mathbb{E}\left[\sum\nolimits_{t=1}^{T}r_{t}(s_{t},a_{t})\mid\pi,P\right],$ (1)
@@ IV-A1 Reduce bi-level coupling to an “oracle” effective loss
$L_{k}^{*}(b_{k})\triangleq\min\nolimits_{\pi^{\text{A}}\in\Pi(b_{k})}\mathbb{E}_{\pi^{\text{A}}}\left[\sum\nolimits_{t=1}^{T}l_{t}(s_{t},a_{t})\right].$ (9)
@@ IV-A2 Dual-based subgradient approximation
Directly optimizing (10) is still challenging because $L_{k}^{*}(b_{k})$ depends on the unknown lower-level MDP dynamics. Our key observation is that, by strong duality of the linear programing (LP) formulation of finite-horizon CMDPs, the optimal dual multiplier $\lambda_{k}^{*}(b_{k})$ of the budget constraint $b_{k}$ provides the sensitivity of the optimal value with respect to the budget. Physically, $\lambda_{k}^{*}(b_{k})$ signal the upper level how much QoS could improve if more RBs were allocated. Thus, $\lambda_{k}^{*}(b_{k})$ constitutes a valid subgradient of loss $L_{k}^{*}(b_{k})$ (see [10] for details). Specifically, for any $b_{k}\in{\mathcal{B}}$, the subgradient of the true loss is given by
@@ IV-A3 Switching cost control via stable OCO updates
While the total episode cost includes the switching term $\alpha\|b_{k}-b_{k-1}\|^{2}$, our update rule operates on (approximate) subgradients of $g_{k}$. As we show in Lemma 2, a decaying step size yields stable budget trajectories with controlled path length, which in turn controls the cumulative switching overhead without requiring a switching regularizer in the update.
@@ IV-A4 Projected subgradient descent with a warm-up phase
    where the step-size is $\eta_{k}=1/[\theta_{g}(k-K_{0}+1)]$.
@@ IV-B1 Empirical estimates and confidence sets
and $L^{\prime}=\log\big(\frac{2SATK}{\delta}\big)$. With probability at least $1-2\delta$, the true model $P$ lies in $\mathcal{P}_{k}$ for all $k$ (see Lemma 1 for details).
@@ V Regret Analysis and Budget Guarantee
In this section, we establish the regret upper bound and cross-level budget constraint satisfaction guarantee for our algorithm. The main technical challenges lie in quantifying the error quantification from the lower-level reinforcement learning under varying threshold to the upper-level optimization updates under coupling and estimated sensitivity feedback. We provide proof sketches for the lemmas and theorems (due to page limits, see complete proofs in our technical report [10]).
@@ V-B Lower-Level Regret Bound
Fix $\delta\in(0,1)$. With probability at least $1-3\delta$, executing BALDE (Algorithm 2) under budget sequence $\{b_{k}\}_{k=1}^{K}$, we have
$\sum\nolimits_{k=1}^{K}\Big(V_{l}(\pi_{k};P)-V_{l}(\pi_{k}^{*}(b_{k});P)\Big)\leq\tilde{\mathcal{O}}\left(ST^{3}\sqrt{AK}/\gamma\right),$
where $\gamma=B_{0}-b_{\mathrm{base}}$.
Lemma 1 provides a high-probability regret guarantee for the lower-level scheduler under an episode-varying budget sequence $\{b_{k}\}_{k=1}^{K}$, which is crucial for coupling with the upper-level provisioning decisions. The bound scales as $\tilde{\mathcal{O}}(\sqrt{K})$, i.e., the average per-episode suboptimality decays as $\tilde{\mathcal{O}}(1/\sqrt{K})$. This $\sqrt{K}$ dependence is optimal and matches the canonical statistical rate for learning under unknown Markov dynamics. The dependence on $\gamma$ makes explicit a fundamental safety-learning tradeoff. $\gamma$ is the minimum safety slack that allows the algorithm to enforce a pessimistically inflated constraint (to guarantee feasibility under uncertainty) while still retaining room for exploration. When $\gamma$ is small, safe learners must behave conservatively, so a blow-up with $1/\gamma$ is unavoidable [14, 13]. Finally, the $T^{3}$ factor is characteristic of finite-horizon optimistic exploration with extended-LP-based planning. It reflects that estimation errors compound along length-$T$ trajectories and that additional factors appear when enforcing constraints pessimistically while maintaining optimism.
@@ V-B2 Step 2 (Safety under a dynamic threshold)
We use a safe baseline policy $\pi_{\mathrm{base}}$ during warm-up phase $K_{0}$ to ensure feasibility when uncertainty is high in early exploration. For $k>K_{0}$, the solver ensures safety by incorporating a cumulative transition bonus $\epsilon_{k}^{\pi_{k}}(P_{k})\triangleq T\mathbb{E}[\sum_{t=1}^{T}{\beta}_{k,t}|\pi_{k},P_{k}]$, we have $V_{d}(\pi_{k};P)\leq V_{d}(\pi_{k};P_{k})+\epsilon_{k}^{\pi_{k}}(P_{k})=V_{\bar{d}}(\pi_{k};P_{k})$. Since the solver ensures $V_{\bar{d}}(\pi_{k};P_{k})\leq b_{k}$, it follows that $V_{d}(\pi_{k};P)\leq b_{k}$ under event $\mathcal{G}$ with a probability $\Pr(\mathcal{G})$.
@@ V-B3 Step 3 (Regret bound)
We decompose the per-episode regret into two terms, i.e., $V_{l}({\pi_{k}};P)-V_{l}({\pi^{*}};P)={(V_{l}({\pi_{k}};P)-V_{\bar{l}}({\pi_{k}};P_{k}))}+{(V_{\bar{l}}({\pi_{k}};P_{k})-V_{l}({\pi^{*}};P))},$ where $V_{\bar{l}}$ is the value under a pessimistic bonus-augmented loss $\bar{l}_{k}=l-mT\bar{\beta}^{p}_{k}$. We show that setting the scaling factor $m\geq T/\gamma$ is sufficient to guarantee $V_{\bar{l}}({\pi_{k}};P_{k})\leq V_{l}({\pi^{*}};P)$. In addition, the first error term $V_{l}({\pi_{k}};P)-V_{\bar{l}}({\pi_{k}};P_{k})$ in the decomposition can be bounded by relating $V_{\bar{l}}$ back to $V_{l}$ via the classic value difference lemma [13, 14, 17] and the cumulative bonus. Finally, summing over $k$ and bounding the cumulative bonus terms $\sum\epsilon_{k}$ by invoking the pigeon-hole principle yields the final regret bound of $\tilde{\mathcal{O}}\left(T^{3}S\sqrt{AK}/\gamma\right)$. ∎
@@ V-C Upper-Level Regret Bound
Lemma 2 isolates two phenomena. First, the term for surrogate regret and switching costs shows that, despite switching costs, the upper-level optimization error scales only as $\tilde{{\mathcal{O}}}(\log K)$ under strong convexity. The key reason switching costs do not destroy the rate is stability, i.e., with $\eta_{k}\propto 1/k$, budget updates shrink over time, so the path length $\sum_{k}\|b_{k}-b_{k-1}\|^{2}$ is bounded by a convergent series. Second, the term for value approximation error captures the penalty for optimizing the surrogate $\hat{g}_{k}$ instead of the true $g_{k}$. This term scales as $\tilde{\mathcal{O}}(\sqrt{K})$ and matches the lower-level learning regret, which confirms that the overall performance is dominated by the difficulty of learning dynamics and the resulting sensitivity.

### v2

@@ Abstract
We study a bi-level online provisioning and scheduling problem motivated by network resource allocation, where provisioning decisions are made at a slow time scale while queue-/state-dependent scheduling is performed at a fast time scale. We model this two-time-scale interaction using an upper-level online convex optimization (OCO) problem and a lower-level constrained Markov decision process (CMDP). Existing OCO typically assumes stateless decisions and thus cannot capture MDP network dynamics such as queue evolution. Meanwhile, CMDP algorithms typically assume a fixed constraint threshold, whereas in provisioning-and-scheduling systems, the threshold varies with online budget decisions. To address these gaps, we study bi-level OCO-CMDP learning under switching costs (budget reprovisioning/system reconfiguration) and cross-level constraints that couple budgets to scheduling decisions. We build an algorithm to solve this bi-level problem by developing several new techniques, including a carefully designed dual feedback that returns the budget multiplier as sensitivity information for the upper-level update and a lower level that solves a budget-adaptive safe exploration problem via an extended occupancy-measure linear program. We establish near-optimal regret and high-probability satisfaction of the cross-level constraints.
@@ I Introduction
Modern networking systems, such as 5G/6G network and edge computing, must support heterogeneous services (e.g., video streaming, augmented reality, virtual reality, and autonomous driving) over shared wireless infrastructure [10]. A central operational challenge in these systems is online resource provisioning under uncertainty. In particular, a network operator must decide how much bandwidth/compute/energy to allocate to a slice over time, balancing quality of service (QoS) against operational expenditure (OpEx). Practical systems exhibit multiple coupled time scales. As illustrated in our motivating example online bandwidth provisioning with queue-aware scheduling (see Fig. 1), the operator provisions an episode-level resource budget at a slow time scale (e.g., bandwidth reserved for a slice over seconds/minutes). This budget directly incurs OpEx and influences the QoS that can be achieved through scheduling. Within each episode, the scheduler performs state-dependent, packet-level actions at a fast time scale (e.g., allocating resource blocks every millisecond). These actions are constrained by the provisioned resource budget and aim to reduce latency and congestion, improving QoS [11, 8]. Moreover, changing the provisioning level across episodes is not free. It induces reconfiguration overheads (e.g., hardware retuning, virtualization/migration cost, and instability), which are naturally captured by switching costs [17, 9]. See Sec. III for details.
Online convex optimization (OCO) is a classical abstraction for sequential resource allocation and has been extensively studied under adversarial and stochastic settings [19, 17, 9]. Recent work incorporates switching costs and related operational constraints to penalize aggressive changes in decisions, which is attractive for network slicing. However, OCO with switching costs typically does not model endogenous Markovian state evolution nor provide safety under unknown state dynamics. For example, the per-round loss depends on the current decision and does not capture the evolution of system states, while in networked systems, state variables such as queue lengths evolve endogenously. Consequently, provisioning decisions that appear cost-effective in the short term may cause persistent congestion and latency violations.
@@
Motivated by the above gaps, we propose a bi-level online optimization and learning framework that integrates online budget provisioning with state-aware safe scheduling (without violating cross-level budget constraints). At the upper level, an OCO operator optimizes and provisions budgets $\{b_{k}\}_{k=1}^{K}$ across episodes, incurring a convex service cost and a quadratic switching cost. At the lower level, the system evolves as an episodic CMDP under varying thresholds. Given the optimized provisioning budget $b_{k}$, the agent must learn a safe policy via RL under unknown dynamics. The two levels are coupled through an episode-wise budget constraint and an episode-wise objective that aggregates cross-level costs.
@@ II-A Upper-Level Budget Provisioning
At the slow time scale, at the beginning of each episode $k\in[K]$, the agent selects an upper-level budget $b_{k}\in\mathcal{B}\triangleq[B_{0},T]$, which can represent the episode-level resource budget provisioned for a network slice (e.g., bandwidth) over seconds/minutes. Choosing $b_{k}$ incurs an episode-dependent convex service cost $f_{k}(b_{k})$, which is not revealed before the decision is made and is revealed at the end of episode $k$. As is typical in online convex optimization with switching costs [9, 6, 17], the service cost $f_{k}(\cdot)$ could change arbitrarily, and it is $\theta_{f}$-strongly convex and differentiable on $\mathcal{B}$ with bounded gradients $\|\nabla f_{k}(\cdot)\|\leq F$.
@@ II-B Lower-Level Stateful Scheduling
At each time $t$ of episode $k$, the agent observes the current state $s_{k,t}$ (e.g., queue backlogs), and then selects a (possibly randomized) scheduling/allocation action $a_{k,t}\sim\pi_{k,t}(\cdot\mid s_{k,t})$, where $\pi_{k}^{\text{A}}=(\pi_{k,t})_{t=1}^{T}$ is a policy with $\pi_{k,t}:\mathcal{S}\to\Delta(\mathcal{A})$. The agent then incurs a stage loss $l_{t}(s_{k,t},a_{k,t})$ (e.g., delay/backlog penalty), consumes resources $d_{t}(s_{k,t},a_{k,t})$ (e.g., used resource blocks), and the system evolves to next state $s_{k,t+1}\sim P_{t}(\cdot\mid s_{k,t},a_{k,t})$ due to unknown arrivals and service dynamics.
@@
$V_{r}(\pi;P)$ $\triangleq\mathbb{E}_{\pi,P}\left[\sum\nolimits_{t=1}^{T}r_{t}(s_{t},a_{t})\right],$ (1)
@@ IV-A1 Reduce bi-level coupling to an “oracle” effective loss
$L_{k}^{*}(b_{k})\triangleq\min\nolimits_{\pi^{\text{A}}\in\Pi(b_{k})}\mathbb{E}_{\pi^{\text{A}},P}\left[\sum\nolimits_{t=1}^{T}l_{t}(s_{t},a_{t})\right].$ (9)
@@ IV-A2 Dual-based subgradient approximation
Directly optimizing (10) is still challenging because $L_{k}^{*}(b_{k})$ depends on the unknown lower-level MDP dynamics. Our key observation is that, by strong duality of the linear programing (LP) formulation of finite-horizon CMDPs, the optimal dual multiplier $\lambda_{k}^{*}(b_{k})$ of the budget constraint $b_{k}$ provides the sensitivity of the optimal value with respect to the budget. Physically, $\lambda_{k}^{*}(b_{k})$ signal the upper level how much QoS could improve if more RBs were allocated. Thus, $\lambda_{k}^{*}(b_{k})$ constitutes a valid subgradient of loss $L_{k}^{*}(b_{k})$ (Please see Appendix B for the complete proof of Lemma 12). Specifically, for any $b_{k}\in{\mathcal{B}}$, the subgradient of the true loss is given by
@@ IV-A3 Switching cost control via stable OCO updates
While the total episode cost includes the switching term $\alpha\|b_{k}-b_{k-1}\|^{2}$, our update rule operates on (approximate) subgradients of $g_{k}$. As we show in Lemma 2, a decaying step size $\eta_{k}\propto 1/k$ yields stable budget trajectories with controlled path length, which in turn controls the cumulative switching costs without requiring a switching regularizer in the update.
@@ IV-A4 Projected subgradient descent with a warm-up phase
    where the step-size is $\eta_{k}=1/[\theta_{g}(k-K_{0})]$.
@@ IV-B1 Empirical estimates and confidence sets
and $L^{\prime}=\log\big(\frac{2SATK}{\delta}\big)$. With probability at least $1-2\delta$, the true model $P$ lies in $\mathcal{P}_{k}$ for all $k$ (Please see Appendix A for the complete proof of Lemma 1).
@@ V Regret Analysis and Budget Guarantee
In this section, we establish the regret upper bound and cross-level budget constraint satisfaction guarantee for our algorithm. The main technical challenges lie in quantifying the error quantification from the lower-level reinforcement learning under varying threshold to the upper-level optimization updates under coupling and estimated sensitivity feedback. We provide proof sketches for the lemmas and theorems (Please see complete proofs in our Appendix).
@@ V-B Lower-Level Regret Bound
Fix $\delta\in(0,1)$. With probability at least $1-3\delta$, executing BALDE (Algorithm 2) under budget sequence $\{b_{k}\}_{k=1}^{K}$, we have ($\gamma=B_{0}-b_{\mathrm{base}}$)
$\sum\nolimits_{k=1}^{K}\Big(V_{l}(\pi_{k};P)-V_{l}(\pi_{k}^{*}(b_{k});P)\Big)\leq\tilde{\mathcal{O}}\left(ST^{3}\sqrt{AK}/\gamma\right).$
Lemma 1 provides a high-probability regret guarantee for the lower-level scheduler under an episode-varying budget sequence $\{b_{k}\}_{k=1}^{K}$, which is crucial for coupling with the upper-level provisioning decisions. The bound scales as $\tilde{\mathcal{O}}(\sqrt{K})$, i.e., the average per-episode suboptimality decays as $\tilde{\mathcal{O}}(1/\sqrt{K})$. This $\sqrt{K}$ dependence is optimal and matches the canonical statistical rate for learning under unknown Markov dynamics. The dependence on $\gamma$ makes explicit a fundamental safety-learning tradeoff. $\gamma$ is the minimum safety slack that allows the algorithm to enforce a pessimistically inflated constraint (to guarantee feasibility under uncertainty) while still retaining room for exploration. When $\gamma$ is small, safe learners must behave conservatively, so a blow-up with $1/\gamma$ is unavoidable [15, 14]. Finally, the $T^{3}$ factor is characteristic of finite-horizon optimistic exploration with extended-LP-based planning. It reflects that estimation errors compound along length-$T$ trajectories and that additional factors appear when enforcing constraints pessimistically while maintaining optimism. Please see Appendix A for the complete proof of Lemma 1.
@@ V-B2 Step 2 (Safety under a dynamic threshold)
We use a safe baseline policy $\pi_{\mathrm{base}}$ during warm-up phase $K_{0}$ to ensure feasibility when uncertainty is high in early exploration. For $k>K_{0}$, the solver ensures safety by incorporating a cumulative transition bonus $\epsilon_{k}^{\pi_{k}}(P_{k})\triangleq T\mathbb{E}_{\pi_{k},P_{k}}[\sum_{t=1}^{T}{\beta}^{p}_{k,t}]$, we have $V_{d}(\pi_{k};P)\leq V_{d}(\pi_{k};P_{k})+\epsilon_{k}^{\pi_{k}}(P_{k})=V_{\bar{d}}(\pi_{k};P_{k})$. Since the solver ensures $V_{\bar{d}}(\pi_{k};P_{k})\leq b_{k}$, it follows that $V_{d}(\pi_{k};P)\leq b_{k}$ under event $\mathcal{G}$ with probability at least $1-3\delta$.
@@ V-B3 Step 3 (Regret bound)
We decompose the per-episode regret into two terms, i.e., $V_{l}({\pi_{k}};P)-V_{l}({\pi^{*}};P)={(V_{l}({\pi_{k}};P)-V_{\bar{l}}({\pi_{k}};P_{k}))}+{(V_{\bar{l}}({\pi_{k}};P_{k})-V_{l}({\pi^{*}};P))},$ where $V_{\bar{l}}$ is the value under a pessimistic bonus-augmented loss $\bar{l}_{k}=l_{k}-mT{\beta}^{p}_{k}$. We show that setting the scaling factor $m=T/(b_{k}-b_{\text{base}})$ is sufficient to guarantee $V_{\bar{l}}({\pi_{k}};P_{k})\leq V_{l}({\pi^{*}};P)$. In addition, the first error term $V_{l}({\pi_{k}};P)-V_{\bar{l}}({\pi_{k}};P_{k})$ in the decomposition can be bounded by relating $V_{\bar{l}}$ back to $V_{l}$ via the classic value difference lemma [14, 15, 18] and the cumulative bonus. Finally, summing over $k$ and bounding the cumulative bonus terms $\sum\epsilon_{k}^{\pi_{k}}$ by invoking the pigeon-hole principle yields the final regret bound of $\tilde{\mathcal{O}}\left(T^{3}S\sqrt{AK}/\gamma\right)$. ∎
@@ V-C Upper-Level Regret Bound
Lemma 2 isolates two phenomena. First, the term for surrogate regret and switching costs shows that, despite switching costs, the upper-level optimization error scales only as $\tilde{{\mathcal{O}}}(\log K)$ under strong convexity. The key reason switching costs do not destroy the rate is stability, i.e., with $\eta_{k}\propto 1/k$, budget updates shrink over time, so the path length $\sum_{k}\|b_{k}-b_{k-1}\|^{2}$ is bounded by a convergent series. Second, the term for value approximation error captures the penalty for optimizing the surrogate $\hat{g}_{k}$ instead of the true $g_{k}$. This term scales as $\tilde{\mathcal{O}}(\sqrt{K})$ and matches the lower-level learning regret, which confirms that the overall performance is dominated by the difficulty of learning dynamics and the resulting sensitivity. Please see Appendix B for the complete proof of Lemma 2.